7b. |5 Compute the work done to move an object from the origin to point (10, 10) along the parabola y = 2 10
7a. [5] Suppose a steady wind blows northeast with w=. The wind exerts a unit of force for each unit of its velocity. Compute the work done to move an object from the origin to point (10, 10) along a straight line.
7b. |5 Compute the work done to move an object from the origin to point (10, 10) along the parabola y = 2 10
Answer #1

Solution:

we have

\vec{\text{F}}=10\sqrt{2}\left \langle \frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}} \right \rangle

r(t) (t, t> ,0 < t < 10

Work Done = / F . dr(t)

/ F (r(t) ) . ґ(t)dt Work Done-

\text{Work Done}=\int_0^{10}\left \langle 10,10 \right \rangle\cdot \left \langle 1,1 \right \rangle\mathrm{d}t

\text{Work Done}=\int_0^{10}20\mathrm{d}t

Work Done 200